A stone is thrown vertically up from the top of a tower with some initial velocity and it arrives on the ground after t1t1 seconds. Now if the same the stone is thrown vertically down from the top of the same tower with the same initial velocity, it arrives on ground after t2t2 seconds. In how much time will the stone take to reach the ground if it is dropped from the same tower?

Grade:11

1 Answers

Arun
6 years ago

Displacement of Stone 1 in time t1 (h) = ut1 – g(t1)2/2
Displacement of Stone 2 in time t2 (h) = ut2 + g(t2)2/2
Displacement of Stone 2 in time t3 (h) = g(t3)2/2

calculate u from the equation 2 and substitute in 1.
calculate h from eq. 3 and substitute in the above derived equation by rearranging and factoring we get (t3)^2 = t1t2

 

Regards,

Arun (AskIITians Forum Expert)